OpenStudy (anonymous):

Two race cars start from the same place and race off in the same direction. The speeds of the two cars are 20 and 40 mph, respectively. How far from the starting point will the two meet if second car left 1 hour after the first car?

OpenStudy (anonymous):

okay let the hours be x

OpenStudy (anonymous):

so one car has 40 mph speed and other 20 so if you want to know after how many hours did they meet it would be 40x = 20(x+1) x+1 because it left one hour earlier

OpenStudy (anonymous):

so you get x = 1 hour

OpenStudy (anonymous):

so the 2nd car is traveling for 1 hour = 40x = 40*1 = 40 miles first car is traveling for x+1 hours,thats 2 hours,so 20*2 = 40 miles explained,so they meet after driving 40 miles